# Divertor shaping with neutral baffling as a solution to the tokamak power exhaust challenge

### v4

@@ 3 Discussion
Using the unique capabilities of MAST-U as a test bed for investigating novel divertor topologies reduces uncertainty in extrapolating current knowledge to reactor class devices by validating both reduced models and exhaust simulations: a crucial milestone for exploring ADCs as a reactor solution. Plasma-molecular interactions, and the critical role they play in the ED and SXD, are one example where model validation has illuminated discrepancies (Figure 5 (h)) that have been reduced with improved rates for $D_{2}+D^{+}\rightarrow D_{2}^{+}+D$. Extrapolating these improved rates to reactors with long-legged, tightly baffled, divertors shows they can make a critical impact on the reactor scale [36]. Additional experiments at higher power (and higher divertor densities) would improve our understanding of how alternative divertors scale with power and are affected by transients, which is an important focus of future MAST-U experiments [56]. Although, the validity of the DLS model is expected to be increased in impurity seeded conditions (see Methods section), the performance of ADCs and the applicability of reduced models requires further validation under impurity seeded conditions, which are important for reactors and requires higher power operation.
@@
ADCs not only improve exhaust performance, requiring lower upstream density/impurity concentration/core radiation on reactors [19], but also improve core-edge compatibility when paired with strong baffling. This enables detached divertor operation without impacting core conditions: a major milestone towards proving the applicability of ADCs in reactors. However, one drawback of some ADCs, including the Super-X divertor, in reactors is that their engineering will be more complex and more costly, due to the additional constraints on space, coils, and magnetic control tolerances [21, 20]. A key finding of our research is that smaller changes in the divertor topology (ED vs SXD) can already strongly improve exhaust performance, as any reactor design needs a compromise between engineering complexity and attractive operating regimes [21, 20]. Moreover, the benefits of the ED are in agreement with reduced model predictions and exhaust simulations. Such reduced models also predict that reducing the X-point radius (particularly on spherical tokamaks) could significantly reduce detachment onset by increasing total flux expansion. This illustrates how our results, indicating that ADCs are a continuum and benefits can already be obtained with more minor shape modifications, pave the way for reactor ADC designs with improved power exhaust and core-edge compatibility, which require less space and have a reduced engineering complexity [21, 20].
@@ 4.2 Divertor detachment, ion source/sink inferences and power balance
Assuming that all volumetric power losses are purely due to hydrogenic radiation as well as dissociation, which is motivated by the observation that hydrogenic radiation estimates from spectroscopic analysis align with the measured total radiation in these conditions [36], the power into the divertor chamber can be estimated by summing $P_{\perp,target}$ and the inferred hydrogenic divertor radiative power loss. This ignores power transfer from the plasma to the neutral cloud through elastic collisions, which can become substantial in the SOLPS-ITER simulations shown (up to 10 % of $P_{SOL}$), in qualitative agreement with the observed rotational $D_{2}$ rotational temperatures [55].
@@ 4.3 Detachment Location Sensitivity (DLS) model
The DLS model formally assumes that all power is dissipated by impurity radiation and that the radiating specie has a constant concentration in the radiating region. This is likely not the case in the MAST-U divertor chamber where the radiative power losses are dominated by hydrogenic interactions [8]. However, the impact of the divertor topology on the detachment onset appears to be more generally applicable outside of impurity radiation dominant conditions.

### v5

@@ 3 Discussion
Using the unique capabilities of MAST-U as a test bed for investigating novel divertor topologies reduces uncertainty in extrapolating current knowledge to reactor class devices by validating both reduced models and exhaust simulations: a crucial milestone for exploring ADCs as a reactor solution. There are, however, key differences between MAST-U and a reactor that must be addressed.
  1. 1.
    Reactors will operate at higher power input and smaller heat flux widths than MAST-U. As a result, heat loads will become so large that impurity seeding will be required to lower target temperatures sufficiently to enable detachment. Understanding how combining total flux expansion, divertor neutral baffling, and poloidal leg length affects impurity-driven power dissipation requires further investigation. Reactor-scale simulations have, however, demonstrated that ADCs enhance power dissipation, enabling operation with target heat fluxes below engineering limits with reduced impurity concentrations [19, 12, 17].
    Although higher power conditions impact neutral transport, plasma-molecular interactions are expected to play a key role in obtaining significant power, momentum and particle dissipation during (deep) detachment according to exhaust simulations that incorporate long-legged, tightly baffled, divertors [36, 12, 42]. The critical role plasma-molecular chemistry plays in the ED and SXD illuminated discrepancies with simulations (Figure 5 (h)) that have been reduced with improved rates for $D_{2}+D^{+}\rightarrow D_{2}^{+}+D$. Extrapolating these improved rates to reactors with long-legged, tightly baffled, divertors shows they can make a critical impact on the reactor scale [36].
  2. 2.
    Reactor-grade operation typically involves H(igh confinement)-mode operation. Current H-mode operation exhibits violent ELMs that result in extremely high heat loads. Although reactors will aim to minimize or suppress ELMs, it remains unclear whether ADCs can effectively mitigate the heat loads of residual (fast, ELM) transients to increase divertor lifetimes.
  3. 3.
    Our MAST-U results show a balanced double-null configuration, with similar particle fluxes reaching the lower and upper outer divertors and negligible power reaching the inner target (Figure 6 [36]). Achieving such up/down balance in reactors poses challenges due to a larger distance between the X-point and poloidal field coils [13] and smaller scrape-off-layer widths ($10-12$ mm for the current MAST-U experiments; $1-2$ mm for STEP [55] and $0.2-0.4$ mm for SPARC [14]). Although the exhaust benefits of double null may be limited according to reactor exhaust simulations [60] and TCV experiments [50], double-null imbalances exacerbates inner target power loading and may necessitate solutions for spherical tokamak reactors, such as a proposed inner target X-Divertor geometry [12, 55] requiring further experimental validation.
    • One concern of outer target optimisation strategies is that the increased outer target connection length will exacerbate the inner target heat load in single null conditions in attached conditions according to reduced models [61]. This was in contrast to detached DEMO Super-X divertor simulations, which indicated inner target heat loads were reduced with outer divertor optimisation [21, 19], requiring further study.
    • This reduced model (for attached conditions) predicts an increase of inner target heat loads by 31 % (ED) and 57 % (SXD) compared to the CD. However, these increased inner target heat loads are reduced when accounting for outer target detachment. Using the ED and SXD outer divertor detachment front as a virtual target, the additional inner target heat loads are reduced below 15 % compared to an outer target attached CD.
Future MAST-U experiments aim to address these key differences and advance towards more reactor-relevant scenarios. Planned upgrades include increased external heating from 4.4 MW to over 10 MW ($>2026$), enabling hotter, more attached divertor conditions. Cryopumping has been installed to reduce divertor neutral pressures and obtain hotter divertor conditions (2025). Advanced scenario development (2025) may enable enable single-to-double-null comparisons. Preliminary results suggest that the benefits of ADCs observed in this study persist in H-mode and may even buffer small ELMs, requiring further investigation [11].
@@
Overall, our results demonstrate that ADCs not only improve exhaust performance, enabling reduced upstream density and likely thus impurity concentration/core radiation in reactors [19], but also improve core-edge compatibility when paired with strong baffling. This allows for detached divertor operation without compromising core conditions: a major milestone towards proving the applicability of ADCs in reactors. However, any reactor design needs a compromise between engineering complexity and attractive operating regimes [21, 20]. The increased engineering complexity of some ADCs, such as the Super-X divertor, remains a significant consideration for reactor designs due to associated cost, space constraints and magnetic control tolerances [21, 20]. Our findings underscore that smaller modifications to divertor topology, such as transitioning from CD to ED, can achieve significant performance gains that are consistent with reduced model predictions and exhaust simulations. ADC design is therefore a continuum - an insight that has implications for reactor designs of DEMO, as well as more compact machines (SPARC, ARC, STEP) [55, 17, 14], and paves the way for designs that optimise power exhaust and core-edge compatibility with reduced engineering and integration demands [21, 20].
@@ 4.2 Divertor detachment, ion source/sink inferences and power balance
Assuming that all volumetric power losses are purely due to hydrogenic radiation as well as dissociation, which is motivated by the observation that hydrogenic radiation estimates from spectroscopic analysis align with the measured total radiation in these conditions [36], the power into the divertor chamber can be estimated by summing $P_{\perp,target}$ and the inferred hydrogenic divertor radiative power loss. This ignores power transfer from the plasma to the neutral cloud through elastic collisions, which can become substantial in the SOLPS-ITER simulations shown (up to 15 % of $P_{SOL}$ [42]), in qualitative agreement with the observed rotational $D_{2}$ rotational temperatures [42].
@@ 4.3 Detachment Location Sensitivity (DLS) model
The DLS model formally assumes that all power is dissipated by impurity radiation and that the radiating specie has a constant concentration in the radiating region. This is likely not the case in the MAST-U divertor chamber where the radiative power losses are dominated by hydrogenic interactions [8] consistent with SOLPS-ITER modelling predictions [37], although impurity radiation could be significant upstream of the divertor chamber entrance [8, 36]. However, the impact of the divertor topology on the detachment onset appears to be more generally applicable outside of impurity radiation dominant conditions.